OpenStudy (anonymous):

MEDAL!!! ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Bob and Susie wash cars for extra money over the summer. Bob's income is determined by f(x) = 6x + 13, where x is the number of hours. Susie's income is g(x) = 4x + 18. If Bob and Susie were to combine their efforts, their income would be h(x) = f(x) + g(x). Assume Bob works 3 hours. Create the function h(x) and indicate if Bob will make more money working alone or by teaming with Susie.

OpenStudy (anonymous):

h(x) = 2x + 5, work alone h(x) = 2x + 5, team with Susie h(x) = 10x + 31, team with Susie h(x) = 10x + 31, work alone

OpenStudy (anonymous):

h(x) = 10x + 31, team with Susie

OpenStudy (anonymous):

Thanks! Think you could help w. one more?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

A function is created to represent the costs of living per person in the family. What restrictions would be made to the domain? The domain would only include integers. The domain would only include positive integers. The domain would only include positive numbers. The domain would include all real numbers.

OpenStudy (anonymous):

The domain would only include positive integers. The integers can represent a person, so it wouldn't be possible to have negative people

OpenStudy (anonymous):

It also can't be positive number only because then how would you explain when one loses money?

OpenStudy (anonymous):

ah, okay I see, thanks. I just was confused with numbers because it was talking about costs, but integers make sense, thanks again:)